### v1

# Fourier Quasicrystals on Preliminary Report

Abstract
This paper has three aims. First, for we construct a family of real-rooted trigonometric polynomial maps whose divisors are Fourier Quasicrystals (FQ). For these divisors include the first nontrivial FQ with positive integer coefficients constructed by Kurasov and Sarnak [41], and for they overlap with Meyer’s curved model sets [59] and two-dimensional [60] and multidimensional [61] crystalline measures. We prove that the divisors are FQ by directly computing their Fourier transforms using a formula derived in [44]. Second, we extend the relationship between real-rootedness and amoebas, derived for by Alon, Cohen and Vinzant [1], to the case The extension uses results in [9] about homology of complements of amoebas of algebraic sets of codimension Third, we prove that the divisors of all uniformly generic real-rooted are FQ. The proof uses the formula relating Grothendieck residues and Newton polytopes derived by Gelfond and Khovanskii [29] . Finally, we note that Olevskii and Ulanovskii [60] have proved that all FQ are divisors of real-rooted trigonometric polynomials for but that the situation for remains unsolved.
2020 Mathematics Classification 52C23; 32A60; 32A27 111This work is supported by the Krasnoyarsk Mathematical Center and financed by the Ministry of Science and Higher Education of the Russian Federation (Agreement No. 075-02-2023-936).
Contents
- 1 Introduction
2 Preliminaries
- 2.1 Notation and Conventions
- 2.2 Tempered Distributions and Radon Measures
- 2.3 Crystalline Measures and Fourier Quasicrystals
- 2.4 Bohr and Besicovitch Almost Periodicity
- 2.5 Bohr and Besicovitch Almost Periodic Multisets
- 2.6 Holomorphic Maps and Multiplicity of their Roots
- 2.7 Representation of Trigonometric Maps
- 2.8 Minkowski Geometry and Mixed Volumes
- 2.9 Genericity for Laurent Maps
- 2.10 Genericity and Uniform Genericity for Trigonometric Maps
- 2.11 Pontryagin Duality for Locally Compact Abelian Groups
- 2.12 Compactification of Groups, Functions, Measures. Multisets
- 3 Examples
- 4 Amoebas and Real-Rootedness
- 5 Sufficient Conditions for Fourier Quasicrystals on
- 6 Conclusions and Future Research
1 Introduction
In 1974 Roger Penrose [67], [32] introduced his first aperiodic planar tilings with fivefold rotational symmetry.
In 1981 de Bruijn [11] described a ”cut and project method” to construct Penrose tilings by projecting a subset of a cubic lattice in five dimensions onto a plane.
In 1983 Paul Steinhardt and his student Dov Levine, inspired by these tilings, proposed a general concept of quasicrystals in a patent disclosure and published it in 1984 [48].
Working independently, in 1984 Dan Schechtman and colleagues [78] synthesized substances exhibiting Bragg diffraction peaks like crystals but whose icosahedral
symmetries were theoretically impossible for crystals. Schechtman was awarded the Chemistry Nobel Prize in 2011 for discovering these physical quasicrystals. In 2019 Steinhardt [80] described the international exhibition he led to Kamchatka in 2011 that found natural physical quasicrystals in meteorites.
The cut and project method to construct models for both aperiodic tilings and physical quasicrystals are similar to the model sets studied a decade earlier by Meyer [54]. Also see [55, 56, 62]. The model sets are discrete but define Radon measures whose
spectrums, defined by their Fourier transforms, are supported on dense subsets. More recent work by Meyer
[57, 59, 60, 61] and Lev and Olevskii [49, 50, 51, 52] focusses on constructing crystalline measures - those whose whose supports and spectrums are both discrete, and a subclass called Fourier quasicrystals.
In [42] Lagarias proved that model sets are
Besicovitch but not Bohr almost periodic. Work
[14, 15, 21, 22, 16, 20, 39] by Favorov and colleagues studied almost periodic sets and their relation to divisors of holomorphic almost periodic functions in tube domains.
In 2000 Kurasov and Sarnak [41] proved that the divisor of every real-rooted univariate trigonometric polynomial is a Fourier quasicrystal and constructed such polynomials from Lee-Yang polynomials. Shortly after in 2000 Olevskii and Ulanovskii [66] proved that every Fourier quasicrystal on is the divisor of a real-rooted univariate trigonometric polynomial. These two results in combination completely characterize Fourier quasicrystals on
As summarized in the abstract this paper has three aims. Achieving these aims required using extensive notation and background knowledge. For this purpose we have included a long section titled Preliminaries. A key result is Fourier quasicrystals with positive integer coefficients are Bohr almost periodic multisets but 3.1 gives examples showing that the converse fails. For these examples the compactifications of multisets are not algebraic varieties. This fact supports our conjecture that the result of Olevskii and Ulanovskii referenced in the abstract holds for though we do not yet have a proof.
We strove to make the paper self contained. However, a detailed explanation of the material in subsection 5.2 would increase the length of this paper to over 100 pages so we merely referenced the seminal paper of Gelfond and Khovanskii [30].
2 Preliminaries
2.1 Notation and Conventions
Throughout this paper we use the following notation and facts.
empty set, composition of maps, homeomorphism, isomorphism.
means ’is defined as’, iff if and only if,
For sets number of elements in
For a locally compact Hausdorff () topological space
space of continuous complex-valued functions on
its Banach subspace with norm
space of uniformly continuous where is a uniform space. e.g. metric space or topological vector space.
its closed subspace of functions that converge to at infinity (for every there exists a compact such that on )
its dense subspace of compactly supported functions,
space of complex-valued Borel (finite) measures on
space of complex-valued Radon measures on
are integer, nonnegative integer, positive integer, rational, real, positive real, and complex numbers.
are the real part, imaginary part, modulus of
Tuples are represented as column vectors.
transpose.
For
maximal number of -independent elements in
Lebesgue Banach space with norm for
If then the group generated by
is isomorphic to
open ball with center radius in
is the Lebesgue measure of the unit ball in
open ball with center radius in
set of by real matrices defining maps
maximum number of -independent rows of
is the general linear group over
is the identity matrix.
and are topological groups under multiplication.
and are compact subgroups.
is the identity and multiplication is denoted by juxtaposition.
is the homomorphism with kernel defined by
induces a homomorphism defined by
thus
and
algebra of Laurent polynomials consisting of sums
where and the spectrum of is finite.
is a frequency of and is its corresponding amplitude.
convex hull of is the Newton polytope of
set of vertices of
is the set of roots of
set of Laurent maps
is the set of roots of
induces a homomorphism
defined by
thus
set of trigonometric polynomials defined by
where and the spectrum of is finite.
is a frequency of and is its corresponding amplitude.
convex hull of is the Newton polytope of
set of vertices of
is the set of roots of
set of trigonometric maps
is the set of roots of
is real-rooted if
If is real-rooted and is discrete then
is the divisor of Here is the point measure at defined in 2.2 and
is the multiplicity of defined in 2.6.
2.2 Tempered Distributions and Radon Measures
For a topological vector space
its dual space consisting of continuous linear functionals
space of Schwartz [76] functions all of whose derivatives exist and decay fast. Its topology is defined by seminorms
space of tempered distributions. Identify by
is the Fourier transform defined by
It is a continuous bijection and
For Fubini’s theorem ([72], p. 164) gives
This allows us to consistently define the bijection by
Let for
is a dense subspace of The
Fourier-Laplace transform of defined by
is the analytic continuation of to an entire function.
Lemma 1
implies there exists depending only on the support of and for every there exists such that
Conversely, (13) implies that
Proof. This is the Paley–Wiener-Schwartz theorem ([34], Theorem 7.3.1).
If then its total variation defined in ([72], p. 116),
is a positive Borel measure, and is a Banach space with norm
([72], Theorem 6.2).
The Riesz representation theorem ([72], Theorem 6.19) implies
that
The point measure at
for every Borel Therefore
where has the inductive limit topology ([81], p.515). A Radon measure is discrete if it
equals
where is discrete (has no limit points)
and Lebesgue measure
on is a Radon measure but not a Borel measure.
The following is from ([3], Definition 2.1).
Definition 1
is a tempered Radon measure if
there exists
such that
Since is dense in tempered Radon measures can be identified with a subclass of tempered distributions. In ([34], Theorem 2.1.7) Hormander proved that all positive distributions are positive measures, hence every positive tempered distribution is a positive tempered Radon measure.
Lemma 2
For positive the following conditions are equivalent.
1.
is a tempered Radon measure.
2.
There exist such that
3.
There exists such that
Favorov ([18], Lemma 1) called property 2 slowly increasing and noted that is equivalent to property 1. The equivalence of properties 2 and 3 follows since
Remark 1
If is Lebesgue measure then the Fourier inversion formula implies
hence The Dirac comb is clearly a Radon measure.
Lemma 3
If is the Dirac comb then or equivalently
which is called the Poisson Summation Formula.
2.3 Crystalline Measures and Fourier Quasicrystals
Definition 2
A Crystalline Measure (CM) on is a discrete tempered Radon measure whose Fourier transform is a discrete Radon measure. A Fourier Quasicrystal (FQ) on is a CM on such that both and are tempered.
Remark 2
Favorov [19] constructed a CM on that is not a FQ.
Lemma 3 implies that the Dirac comb is a FQ. For and the affine transformation on gives am affine transformation on discrete Radon measures If is a CM with then a computation gives
so is a CM. Furthermore if is a FQ then is a FQ. Clearly CM and FQ are closed under finite linear combinations. Furthermore, if is a CM (resp. FQ) on and is a CM (resp. FQ) on then their product is a CM; FQ on
Definition 3
A CM on is called trivial if it is a finite linear combination of affine transformations of the Dirac comb. This occurs iff its support is a finite union of arithmetic sequences. All trivial CMs on are FQ’s. A FQ on for is called trivial if it is a finite linear combination of affine transformations of products of FQs.
2.4 Bohr and Besicovitch Almost Periodicity
The theory of uniform almost periodic functions on was pioneered in three papers [6] by Harald Bohr and extended to mean-squared almost periodic functions in [5] by Abram Besicovitch. We summarize their results extended to define almost periodic measures and relate them to tempered distributions and Fourier quasicrystals in Propositions 1 and 2.
Definition 4
is relatively dense if there exists such that ([7], p. 32, Main Definition) is almost periodic if for every the set
is relatively dense. is called the set of -almost periods of
We call these Bohr almost periodic functions, denote the set of them by and record that
1.
([7], Theorem I)
2.
([7],Theorem II)
3.
([7],Theorem III) closed under addition and hence
4.
([7],Theorem IV) closed under multiplication
5.
([7],Theorem V) closed under uniform closure.
- 6.
- 7.
8.
([7], p. 60-65) This uniqueness is equivalent to Parseval’s equation
(19) -
9.
([7], p. 80-88, Fundamental Theorem) uniform closure of
(19) implies that the spectrum
is countable. For we call the set of translates of its orbit and the uniform closure its orbit closure. In [8] Bochner proved
In [5] Besicovitch defined the generalized almost periodic functions completion of with respect to the norm
Clearly In [42] Lagarias proved that functions associated with physical quasicrystals and Meyer’s model sets are in This result is explained in detail in [56] by Meyer.
Remark 3
A subset of an abelian topological group is syndetic if there exists a compact such that For syndetic means relatively dense. These concept play a central role in topological dynamics where its relation to minimal systems was pioneered in ([31], Chapter 4) by Gottschalk and Hedlund. For and define the translation of by and the orbit of , Then is a uniformly equicontinuous group of transformations of ([31], 11.08) so for every and every extends to a map ([31], 4.44) implies that iff is compact. This gives Bochner’s [8] characterization of almost periodicity: is Bohr almost periodic if every sequence of translations of has a subsequence that converges uniformly. Moreover ([31], 4.48) implies that the following conditions are equivalent:
1.
2.
is a compact minimal closed subset of invariant under
3.
there is a unique group structure on which makes it a compact topological group and the map defined by has a dense image.
If define and by for Then The Stone-Weierstrass theorem ([71], A14) implies that is dense in so can be uniformly approximated by hence can be uniformly approximated by Since Bohr’s fundamental theorem follows. The preceding results illustrate the power of abstraction embodied in topological dynamics
If and
then
the right side of (24) converges uniformly independent of the ordering of terms.
([7], p. 89, Convergence Theorem) shows that this always occurs whenever
is linearly independent over
If (23) fails then
the right side of (24) must be computed as a limit of Fejer sums as explained in ([7], p. 66).
The convolution of and
is in and
Definition 5
A Radon measure is a Bohr (resp. Besicovitch) almost periodic measure if for every the convolution is a Bohr (resp. Besicovitch) almost periodic function. A Bohr (resp. Besicovitch) almost periodic distribution is a tempered Radon measure such that for every is a Bohr (resp. Besicovitch) almost periodic function.
If is a Bohr almost periodic measure define by
where and (25) ensures that is well defined. The spectrum
is countable since there exist such that for all
Lemma 4
If is a Bohr almost periodic measure then iff
Then
is called an atomic measure.
Proof. Meyer proved this ([58], Theorem 5.5).
Remark 4
Favorov ([17], Lemma 1) proved that if is a temperate measure, is a point measure, and for some then is an almost periodic distribution and for every converges absolutely.
We record the following observation without proof.
Remark 5
If is a Bohr almost periodic measure and (28) fails. Then if there exists such that and is differentiable and has derivative except on where has a jump equal to at every If then the description of is more complicated.
Definition 6
A Radon measure on is translation bounded if there exists such that
Since is contained in a union of balls of the form Lemma 2 implies that every translation bounded Radon measure is a tempered Radon measure.
Proposition 1
If then is a Bohr almost periodic measure iff is translation bounded and is a Bohr almost periodic distribution.
Proof. Let be a Bohr almost periodic measure. Let be the Banach space of consisting of whose supports are contained in and let For every define by
Clearly each is bounded on since
Furthermore
since implies it is bounded. Therefore the Banach-Steinhaus theorem ([72], 5.8) implies that there exists such that
This shows that is translation bounded. Let We conclude the proof of the only if part by proving that Let be a sequence such that equals on and for It is easy to show that is a bounded continuous function and converges uniformly to Since each and is closed under uniform limits,
To prove the if part assume that is a translation bounded measure and a Bohr almost periodic distribution. Let We conclude the if part of the proof by showing that
Let such that Then is a bounded continuous function and is the uniform limit of hence
Sergey Favorov sent us the following result in a personal communication.
Proposition 2
Every positive FQ on is translation bounded and is a Bohr almost periodic distribution hence is a Bohr almost periodic measure.
Proof. Let be a positive FQ so where is discrete, and is a tempered distribution. Therefore there exists such that
Let Then where hence
Since there exists such that for Therefore there exist with
This proves and has an absolutely converging Fourier-Bohr series so is a Bohr almost periodic distribution. Choose values in whose support is in and equals on Since it is bounded and positive implies hence
so is translation bounded. Proposition 1 then implies that is a Bohr almost periodic measure.
2.5 Bohr and Besicovitch Almost Periodic Multisets
In [40] Krein and Levin introduced almost periodic multisets in in their study of the zeros of holomorphic almost periodic functions. In [21] Favorov and Kolbasina studied almost periodic multisets in We discuss almost periodic multisets in
Definition 7
A multiset in is a pair where is discrete and We let denote the set of multisets in and denote its subset where on
can be represented by a sequence where for every for values of Two sequences and represent the same multiset iff there exists a permutation
such that
Therefore multisets are identified with equivalence classes of sequences with values in
Define the distance between multisets and
where is represented by a sequence is represented by a sequence and ranges over all permutations of We observe that the metric space is complete since every Cauchy sequence converges.
Remark 6
The restriction of to is the Hausdorff distance explained in ([63], p. 279, Exercise 7)
For define the translation by
Since these translation are isometries the dynamical system is uniformly equicontinuous. Following the logic used to analyze Bohr almost periodic functions in 2.4, we define to be Bohr almost periodic if for every the set of -almost periods
is syndetic. Therefore is Bohr almost periodic iff its orbit closure is compact and then the orbit closure admits the structure of a compact group with as the identity. Define the sum of multisets and by
where on on and on Associate to a multiplicity set the measure
We record the following observation without proof.
Example 1
If and is irrational, where characteristic function of where characteristic function of and then is Bohr almost periodic,
and the group is a trivial FQ.
In [21, 39] Kolbasina and Favorov proved that a multiset in is Bohr almost periodic iff is almost periodic and then its density, defined by
exists and convergence of the sum is uniform in We define a multiset in to be Besicovitch almost periodic if is a Besicovitch almost periodic measure. In [16] Favorov characterized Besicovitch almost periodic multisets in by their translation properties.
Definition 8
A multiset is uniformly discrete if on and where and are distinct elements in is relatively uniformly discrete if it equals the sum of a finite number of uniformly discrete multisets.
For a multiset define its index by
We observe that is relatively uniformly discrete iff and then equals the sum of uniformly discrete multisets. All Bohr almost periodic multisets in are relatively uniformly discrete.
Remark 7
Assume that is uniformly discrete and define where are distinct elements in If is Bohr almost periodic and then the group structure of is isomorphic to If and then there is a unique bijection such that This defines a covering fiber bundle over ([44], Theorem 5) shows that the toral compactification of is a submanifold of homeomorphic to and the homotopy of the embedding is injective. We conjecture that this homotopy determines the Galois structure of the covering. Furthermore if is only relatively uniformly discrete then this construction gives a more complicated branched covering and we conjecture that its Galois structure is determined by the homotopy of the toral compactification of which equals the image in of a manifold homeomorphic to
2.6 Holomorphic Maps and Multiplicity of their Roots
A domain is a nonempty open connected set. A map is holomorphic if each component is holomorphic. is isolated if there exists such that and Assume that is an isolated root of For and the multiplicity of equals its logarithmic integral
where is the one-cycle consisting of a circle with center at bounding a closed disc contained in and containing only the root and oriented so is positive (counterclockwise orientation). For the multiplicity can be computed by a Bochner-Martinelli integral over a sufficiently small -dimensional sphere centered at ([82], p. 16, Eqn. 1), ([77], p. 286, Eqn. 3). Alternatively, its multiplicity can be computed by an integral
over the -cycle
where is a neighborhood of with are sufficiently small to ensure that and is oriented so that the differential form We call a Grothendieck cycle since (44) is the Grothendieck residue of the Jacobian ([82], p. 16). The Bochner-Martinelli integral representation of multiplicity gives:
Lemma 5
(Rouche) If is a bounded domain with Jordan boundary continuous, holomorphic, and then has as many zeros in as counting multiplicity.
Proof. Follows from ([77], Theorem 2, p. 288).
Definition 9
A (horizontal) tube domain is a subset where is nonempty, open and bounded.
The set of bounded holomorphic functions on is a Banach space with norm If the set of roots of is discrete we let denote the multiset where for every We also let denote the associated Radon measure The interpretation will be clear from context.
Definition 10
A holomorphic map is Bohr almost periodic if it is bounded for every there exists a syndetic such that
denotes the set of holomorphic almost periodic maps For the remainder of this subsection assume that is a tube domain, and is discrete. Define the lower density of by
For let Rouche’s lemma 5 gives:
Corollary 1
If then
Proof. Let Choose such that and Then has roots in counting multiplicity. Define Then hence there exists a syndetic with
Then for hence Rouche’s lemma 5 implies that has roots in
counting multiplicity. Since
is syndetic it must contain a subset with lower density
such that such that
for every distinct
Then the open sets are pairwise disjoint and each has roots of counting multiplicity. Therefore
which concludes the proof.
For define the compact set and the closed set
Recall that is the uniform closure of the set of translates of by
Corollary 2
If is discrete, then for every
Proof. Assume to the contrary that Then there exists a sequence with (21) implies is compact and bounded implies is compact. Therefore we may replace with a subsequence such that for some and for some Then
Since is discrete there exists with and For sufficiently large and Rouche’s lemma 5 implies has a root hence But since and This contradiction concludes the proof.
Corollary 3
If there exists such that is discrete and then is a Bohr almost periodic multiset and
Proof. A similar argument used in the proof of Corollary 1 shows that there exists and such that each connected component of is the union of at most sets Let be the sequence interiors of these connected components. Corollary 2 gives
There exists a syndetic with
Therefore
This implies that which implies that is Bohr almost periodic (considered as a multiset, or equivalently, as a measure). The second assertion follows since the metric on induced by is uniformly continuous with respect to the metric induced by
2.7 Representation of Trigonometric Maps
For we write if the exists such that This is an equivalence relation. We say and are in the same class.
Definition 11
is minimal if
Proposition 3
If and then
and
If and then there exists a minimal and such that For every minimal and that satisfies there exists such that and
Proof. The first assertion follows from a direct computation. Let and whose columns form a -basis for and For every there exists a unique such that Then for every entry of (LABEL:hatWell) implies that
where is defined by
Therefore where Since Since the column of are -independent, so is minimal. To prove the third assertion note that hence there exist with and and Then Since the columns of are -independent, hence hence Let Then (LABEL:rhomhatN) implies
Lemma 1 implies that has a dense image hence so
Corollary 4
If and is a tube domain then
Proof. For and hence
Proposition 3 implies there exists and such that Let is a component of and the corresponding component of Then Let For the triangle inequality gives
where
(51) gives
acts on to give a dynamical system
This action is uniformly equicontinuous and since every orbit is almost periodic so for every there exists a syndetic such that
For every choose
(52), (54) and the fact that implies gives
2.8 Minkowski Geometry and Mixed Volumes
For any compact convex –dimensional volume of The Minkowski sum of compact convex and is In [4] Bernshtein defined the mixed volume of an -tuple of compact convex sets
and noted that it differs from the classical definition by the factor We will use Bernshtein’s definition in the remainder of this paper.
Lemma 6
iff there exists points with is linearly independent. Then
Proof. ([75], Theorem 1.5.7, Theorem 5.1.8).
Lemma 7
If the vertices of are in then
Proof. ([13], Theorem 3.9).
2.9 Genericity for Laurent Maps
The fundamental theorem of algebra ensures that if then
has roots in counting multiplicity. The situation for Laurent maps is more complicated even for a system of two linear equations in two variables. Consider where
Let
Then where is the triangle with vertices
A computation shows iff
and then is called generic.
Otherwise or is unbounded and has no isolated points. We note that each of the three inequalities that characterize genericity involves amplitudes associated with vertices on one side of
For is analytic so is finite iff it is compact
([10], Proposition 1, p. 31) and genericity of was
defined by Bernshtein [4] who proved
Lemma 8
If is generic then is finite and
If is not generic then is empty or noncompact,
and
Here is the set of isolated roots of
Fix and for let be a polytope whose set of vertices
The set of with
can be parameterized by the set where and
We define genericity for Laurent maps.
For and
define
Definition 12
is generic if for every
is a face of orthogonal to with dimension so there are only a finite number of and Bernshtein noted each
”depends essentially on smaller number of variables” and therefore
”does not have any zeros in the general case.” He shows that the set of non generic are parameterized by a proper algebraic subset of
Therefore if is generic and and the amplitudes of are sufficiently close to the amplitudes of then is generic.
We illustrate genericity with an example more complex that the previous example. Consider the set of
where
and
is the square with vertices These set of Laurent polynomials are parametrized by
Then
where all coefficients are in
Let Then
is generic when all Clearly the last four Laurent maps above have no roots. So a direct computation shows that fails to be generic on the algebraic set defined by the union of four hypersurfaces defined by
If is generic then a calculation shows that has two roots counted with multiplicity and
2.10 Genericity and Uniform Genericity for Trigonometric Maps
Every nonzero has the form
where and
are real. Then for sufficiently large
dominates the other terms in and for
sufficiently large dominates the other terms in Therefore is bounded. Moreover, is uniformly bounded for sufficiently close to and sufficiently close to Since is a nonzero univariate holomorphic function, is empty iff otherwise it is an infinite discrete set. Corollary 4 implies is Bohr almost periodic. Hence corollary 3
implies is Bohr almost periodic so (41) implies that
its density exists. A contour integration argument gives
The situation for trigonometric maps is much more complicated and we must impose stringent conditions on
to ensure that is bounded, is discrete and its density exists, and that these properties are preserved under sufficiently small perturbations.
We define genericity for trigonometric maps.
For and
define
Definition 13
is generic if for every
Let and such that For every define
Hence Then is the uniform closure of the set of translates of by elements in We note that and only the arguments but not the moduli of the amplitudes of differ from those of
Definition 14
is uniformly generic if is generic for all
Gelfond [27] announced that if is uniformly generic then is discrete, is bounded, and there exists with
This implies that the density
Remark 9
Gelfond deals with holomorphic Bohr almost periodic maps on vertical tube domains and counts roots whose imaginary parts lie in a cube He calls his condition properly unfolded. We avoid this term here because we reserve it for a condition that for every consists of a single vertex for some This condition implies uniform genericity but is much stronger. Gelfond’s result applies to trigonometric maps since corollary 3 implies they are Bohr almost periodic. For trigonometric maps the indicator diagrams of the components are their Newton polytopes so Gelfond’s condition ([27], System 3, p. 155) applies and the requirement that be uniformly generic is stated in ([27], (c), p. 156) ”it is required that this closure does not contain consistent systems” meaning with Gelfond’s paper does not provide a proof but references his preprint [26]. He also references a paper of Hovanskii [36], in which 4 out of 7 references are preprints, and a paper by Kazarnovskii [37] in Russian. Kazarnovskii’s paper [38] derived asymptotic distribution of roots for exponential sums whose exponents are not necessarily purely imaginary and thus not trigonometric polynomials.
Proposition 4
If is uniformly generic then there exists a tube domain such that for all sufficiently close to and is a Bohr almost periodic multiset contained Moreover, is real-rooted iff the density of real roots equals the density of all roots.
Proof. Follows from corollaries 1, 3, and 4.
Levin ([47], p. 268, Lemma 1) derived the following result for
Proposition 5
If is uniformly generic and is a tube domain containing is the set of such that whenever or Then
Proof. Follows from corollaries 2 and 4.
2.11 Pontryagin Duality for Locally Compact Abelian Groups
Let be a locally compact abelian group. A continuous homomorphism is called a character of The Pontryagin dual of is the set of characters (continuous homomorphisms) equipped with multiplication and the topology of uniform convergence on compact subsets. is a locally compact abelian group with identity the constant function and is discrete (resp. compact) iff is compact (resp. discrete). If is the additive group of a field then is isomorphic to though not in a unique way. is connected iff is torsion free (no elements of finite order). If and are locally compact topological groups and is a continuous homomorphism then its Pontryagin dual defined by
is a continuous homomorphism. Moreover, is injective (resp. has dense image) iff has dense image (resp. is injective). Pontryagin’s duality theorem ([71], Theorem 1.7.2) implies that the inclusion defined by
is a isomorphism of topological groups, so we identify with and if we identity with Therefore characters on separate points of and iff If is a closed subgroup of a locally compact abelian group define
Pontryagin duality gives If is a closed subgroup of then the quotient is a locally compact abelian group and the map defined by is a continuous surjective homomorphism with kernel Hence iff iff there exists iff with Therefore if is a closed subgroup of then iff there exists such that and if is a closed subgroup of then iff there exists such that
Lemma 9
If and then
Proof. The first assertion follows since iff there exists such that iff iff The second since The third since implies and then iff there exists such that iff iff
2.12 Compactification of Groups, Functions, Measures. Multisets
Definition 15
A compactification of a locally compact abelian group is a pair where is a compact group and is a continuous homomorphism with a dense image. Clearly is abelian and if is connected then is connected. The Bohr compactification is when where equals with the discrete topology and and is the identity homomorphism.
The Bohr compactification has the following universal property: If is any compactification of then there exists a surjective homomorphism such that If is a compactification of we define its spectrum If is the group generated by with the discrete topology then Pontryagin duality gives Let the algebra of finite linear combinations of characters on Since separates points and is closed under complex conjugation, the Stone-Weierstrass theorem ([71], A14) implies that it is dense in For let the discrete subgroup of generated by
Lemma 10
If and then and where integration is with respect to Haar measure on Let be the inclusion map. If then where and is a compactification of Furthermore and there exists a unique such that
Proof. Since is dense in can be uniformly approximated hence can be approximated by so The integral identity occurs because if else Since is discrete is compact and since is injective has a dense image hence is a compactification of Pontryagin duality implies that and hence Then so there exist a sequence with and Since there exists a sequence such that Since has dense image Since converges it is a Cauchy sequence hence is a Cauchy sequence so converges to a function Then
Definition 16
Let denote the set of Borel measures on a compact abelian group The Riesz Representation theorem ([71], Theorem 6.19) implies that is the space of continuous linear functionals on For we define its Fourier transform by
and its spectrum
Definition 17
An approximation of the identity on is a sequence that is nonnegative, and for every open set containing there exists such that support Then converges weakly to
Proposition 6
If is a Bohr almost periodic measure on then there exists a compactification of and such that
The measure is called the trace of with respect to and it is unique.
Proof. Let be the discrete subgroup of generated by then construct the compactification of as in the proof of Lemma 10. Let be an approximation of the identity on Since is translation bounded
Since is a Bohr almost periodic measure there exists a sequence of such that Furthermore are nonnegative and so Therefore the Banach-Alaoglu Theorem ([73], 3.15) implies that there exists a subsequence of that converges weakly to a measure on Replacing the sequences with this subsequence and applying (25) and Lemma 10 gives
Since is dense in is uniquely determined
Remark 10
The trace of with respect to the Bohr compactification equals the image of under the Bohr map ([2], Section 7).
3 Examples
Subsection 1 constructs a family of pairs of Besicovitch almost periodic subsets of whose sums are Bohr almost periodic sets. Their construction is based on their toral compactifications described in subsection 2.10 and ([44], Section 3). Subsection 2 constructs, for every a parametrized family of real-rooted trigonometric maps whose divisors are Fourier quasicrystals.
3.1 Almost Periodic Multisets
We parametrize by with coordinates Let with irrational and define the homomorphism by
and define the differential -form on
Let be irrational with and define -chains oriented in the direction of increasing coordinate.
The compactification of gives a foliation of whose leaves are the cosets of and the constraint on ensures that and are transverse to the foliation. For define the multiset where
and its associated measure are Meyer model multisets [62], ([44], Definition 2) and therefore Besicovitch but not Bohr almost periodic. The compactifications of defined in ([44], Definition 6), is The compactification of was defined in ([44], Definition 6) by a limit
and expressed in ([44], Remark 4) by the integral
Proposition 6 then gives
Since never vanishes A computation gives that for every interval with
iff which proves that the Fourier-Bohr expansion of
does not converge absolutely for
is uniformly discrete, is a positive Radon measure , and and hence are tempered distributions
but is not a Radon measure so is not even defined.
The sum is a Bohr almost periodic multiset because is a closed curve in transverse to the foliation given by
Letting we compute
which is clearly the density of
3.2 Fourier Quasicrystals
Example 2
Fix Let
chosen to make uniformly generic (as will be described below)
where are -independent, and
with
For define by
and by
Lemma 9 implies that is injective and is a compactification of
Lemma 11
is injective, and is real-rooted.
Proof. (76) implies
Since there exists such that therefore hence is uniquely determined by so is injective. implies and equality follows since both sets are irreducible algebraic subsets of of codimension If then
so all have the same sign as . But if for then
so all have the opposite sign of and
so
has the opposite sign of
and is real-rooted.
Define Then and
since is real-rooted
and
Since is transverse to
so the roots of have multiplicity
and
We now compute
is the subgroup
so is a subgroup of
A computation gives iff there exists such that
Expressing this in matrix form gives
Since there exists such that Expressing in block form
where gives
Therefore
and since all roots have multiplicity
Since is transverse to in each root of is transformed into a unique root of and by the implicit function theorem is an analytic function of Therefore the chain rule gives
This equation determines, for each a real-analytic function and the set is bounded. This fact implies that so then (83) gives
A subset of of the form where and is called an infinite arithmetic progression.
Lemma 12
does not contain an infinite arithmetic progression.
Since are rationally independent is injective. Assume to the contrary that contains an arithmetic progression Then is contained in a finite rank subgroup of Define Then and is a finite rank subgroup of The solution of Lang’s conjecture ([12], Theorem 10.10.1) implies is contained in the finite union of translates of torus subgroups of of dimension Since does not contain the translate of any torus subgroup of dimension and hence is finite. Since is injective is finite. This contradiction concludes the proof.
Remark 11
In [43] Lang’s conjectured that if is a finite rank subgroup of and is a proper variety then is contained in a finite number of cosets in of subgroups of isomorphic to for In [53] proved this for and in [45] Laurent proved this for If then a result in [70] of Ritt implies that if contains the arithmetic progression then there exists such that
Let The trace with respect to the compactification derived in Proposition 6 equals
where is a -cycle oriented by the counterclockwise orientation of and is the differential -form
where and Therefore
Theorem 1
is a Fourier quasicrystal.
Proof. and
(88) implies hence it suffices to prove that there exists a polynomial with
Since and for are power series in nonnegative powers of whenever Since is real, whenever If then
Let If then either all or all and Then satisfies the inequality in (89). The proof is finished.
Remark 12
For
is equivalent (under the automorphism ) to the Laurent polynomial that Kurasov and Sarnak used ([41], Equation 41) to construct the first nontrivial FQ.
4 Amoebas and Real-Rootedness
4.1 Background
In 1952 in their study of phase transitions in statistical physics Lee and Yang introduced a family of and proved ([46], Appendix II) that
so has all its zeros in the unit circle. This family is
where is symmetric with and is summed over all subsets of
Subsequently more general polynomials satisfying (91) were studied and named Lee-Yang polynomials.
In 2010 Ruelle ([74], Theorem 3) completely characterized them.
In 2005 Passare and Tsikh interpreted (91) in terms of the amoeba notion and used amoeba properties to prove ([69], Theorem 2) that polynomials defined by (92) satisfy (91) . To explain the amoeba notion define the logarithmic moment map
by
is a homomorphism from onto with kernel A subset of is called an algebraic variety if it is the set of common zeros of a set of Laurent polynomials. By the Noetherian property of the ring of Laurent polynomials, this set of polynomials can be chosen to be finite. is called a hypersurface if the set has one element and then has codimension is called a complete intersection if it equals the intersection of hypersurfaces and has codimension The amoeba of an algebraic variety is
In ([25], p. 194, Definition 1.3) Gelfand, Kapranov and Zelevinsky defined amoebas for hypersurfaces, in ([25], p. 195, Corollary 1.6) they proved that the complement of the amoeba of a hypersurface is a finite union of open convex subsets. Their work was extended in [23, 24, 68]. Clearly is a Lee-Yang polynomial iff or equivalently
In [1] Alon, Cohen and Vinzant used amoebas to prove that every real-rooted admits the representation
where
is a Lee-Yang polynomial, and
We observe that (95) holds if is replaced by a matric sufficiently close to a condition we call -stable.
This section proves Theorem 2, which represents uniformly generic real-rooted trigonometric maps by -stable Laurent maps. Here
This concept generalizes the concept of Lee-Yang polynomials.
4.2 -stable Laurent Maps
Definition 18
For a rank matrix is -stable if
for all in some open neighborhood of
Lemma 13
If have rank and is -stable then for some stable For we can choose and to be a Lee-Yang Laurent polynomial.
Proof. acts on the Grassmann manifold consisting of -dimensional subspaces of An element in is rational if it has the form where has rank The existence of the Smith normal form [64, 79] for integer matrices implies that acts transitively on set of rational elements in Since the set of rational elements is dense, every orbit is dense. The hypothesis implies that there exists an open such that and
Since every orbit is dense, there exists such that so is open and Define Then
Since for all in some open neighborhood of is -stable. This proves the first assertion. To prove the second assertion assume that is nonzero and that a Laurent polynomial is -stable. Then there exists an open such that and
Then there exists with such that The Smith normal form implies that there exists whose first column equals By adding sufficiently large positive integer multiples of to the last columns of we construct such that
Define Then for every
so is a Lee-Yang Laurent polynomial.
4.3 Proof of Proposition 7
The following result is an extension to Laurent maps of the result for Laurent polynomials by Alon, Cohen and Vinzant ([1], Corollary 3.3).
Proposition 7
If and has rank and is uniformly generic, then
Proof. Define
Since is a variety and it suffices to prove that
If then hence
so
Since is closed,
Proving the reverse inclusion is more challenging. We use proof by contradiction. Assume that there exists
Since is closed there exists such that
therefore does not vanish on the set
Proposition 4 implies rank hence is bounded. Since is invariant under translation by elements in is a tube domain and Corollary 4 implies that Moreover the Hausdorff distance
Proposition 5 implies there exists such that
(106) and the fact that is surjective implies there exists such that
(106) implies there exists such that
and (112) and (113) implies therefore (111) implies
For define by
(111), (112), (113) and (115) imply
(104) and (114) imply there exists a sequence such that
Therefore
and the convergence is uniform over Choose sufficiently large that
so has no roots in This contradicts (116) and concludes the proof.
4.4 Proof of Theorem 2
Theorem 2
If is uniformly generic and real-rooted and then there exists a rank matrix with and an -stable with
Remark 13
The proof of Theorem 2 requires the following result.
Lemma 14
If and rank then there exists such that
for sufficiently close to
Proof. Let and let be the boundary of and The hypothesis implies that for every the open ray
Since each is a finite union of open convex subsets there must exist and such that the line segment is contained in one of these open convex subsets. Therefore there exists an open cone such that
and
Since is an open cover of there exists a finite collection whose union contains Define and Then
hence if is sufficiently close to then
The proof of Lemma 14 is concluded since
We complete the proof of Theorem 2 by contradiction. Assume that is uniformly generic and real-rooted and let be the representation in Proposition 3. Since is generic, rank Assume to the contrary that is not -stable. Then Lemma 14 implies that for every neigborhood of there exists and We may assume that is sufficiently small is uniformly generic so there exists with
Proposition 7 then implies
Let and be spheres radii and in oriented in a positive direction to define cycles in and in Then the homology classes in because the point lies between the spheres. However, these cycles are homotopic in as the following argument shows. Define
Then and define homotopies that deforms the -cycles and to homologous cycles in Therefore the cycles and are homologous in so in This means that the homology inclusion
is not injective. This violates ([9], Theorem 1) and concludes the proof.
5 Sufficient Conditions for Fourier Quasicrystals on
This section proves the following result:
Theorem 3
If is real-rooted and uniformly generic then is a Fourier Quasicrystal on
5.1 Approximation Strategy
Since is dense in it suffices to prove the existence of a discrete and such that both and are tempered Radon measures and
Assume that is real-rooted and generic and Theorem 2 implies there exists a rank matrix and an -stable Laurent map such that Define a sequence
where converges to Since is -stable we may assume that each is real-rooted and generic.
Lemma 15
Proof. Since each component of (resp. equals (resp. ) times the corresponding component of where convergence componentwise with respect to the Hausdorff metric. The conclusion then follows from Lemma LABEL:gelfond.
Proposition 8
Proof. For let be the characteristic function of (11) implies
Let be a tube domain containing and are Bohr almost periodic in and converges to uniformly on compact subsets of so Rouche’s Lemma 5 implies that converges Therefore the first term on the right above converges to implies Therefore the second term on the right above
converges to as since has fast decay. Similarly the third term on the right above converges to
as
Express
where
is the least common multiple of the denominators of the entries in
and
(3) gives
where
Since and hence is generic Lemma 8 implies is finite. Then since is real-rooted so there exists a finite with
For define
If then and the multiplicities satisfy
since and multiplicity is preserved under the holomorphic local homeomorphisms. Therefore
The Poisson summation formula Lemma 3 and its consequent (15) imply
Definition 19
For and define by
Define differential -forms on
where and
Define -chains in
where is the Grothendieck cycle defined in (45) with replaced by
Lemma 16
5.2 Grothendieck Residues and Toric Geometry
In ([30], 1.2) Gelfond and Khovanskii introduced the following concept:
Definition 20
Let be an -tuple of polytopes in Then each face of their Minkowski sum is a Minkowski sum where each is a unique face of called a summand of The tuple is called unfolded if every face of has at least one vertex among it summands. Otherwise it is called folded.
The Laurent maps in in 2.7 have Newton polytope tuples where and are identical squares. Therefore is a square each of whose sides has summands consisting of sides not vertices. Therefore is folded. We record the following observation without proof. is unfolded iff for every there exists and a vertex such that
It is easy to show that if and the components of
satisfy (143) then is generic.
If a tuple of polytopes is unfolded then to each vertex of their sum is associated an integer called its combinatorial coefficient, see [28, 29, 30].
If is a trigonometric map and is unfolded, then for sufficiently close to (with respect to its frequencies and corresponding amplitudes) is unfolded
and whenever is a vertex of the sum of the polytopes in
and is the vertex of the sum of the polytopes in
that is closest to
We may assume that each is unfolded so is unfolded since its entries differ from those in by a constant multiple.
For every define the Newton cycle oriented so
is positive.
Proposition 9
If is unfolded and then for every there exists such that
1.
2.
In the restriction of to the monomial corresponding to dominates the other monomials in so that for some
(144) such that for all
3.
is homologous to the Newton -chain defined by
(145) where is the combinatorial coefficient for the vertex
4.
Then Lemma (16) and Stokes theorem imply
Proof. These results are derived in ([30].
5.3 Proof of Proposition 10
Proposition 10
If is real-rooted and is unfolded, then is a Fourier Quasicrystal on
The proof of this will occupy the remainder of this subsection. Define Then is a rational function holomorphic in A computation gives
therefore
where is the constant term in the Laurent expansion of the restriction
Express (149) in terms of the components of where for The chain rule gives the trigonometric series
when or equivalently is the constant term in this series.
Lemma 17
There exist and such that and are tempered Radon measures and
Proof. Since is a linear combination of and the coefficient of is the coefficient of in the trigonometric series expansion of and The frequencies appearing in this series are in the subgroup generated by This subgroup equals (149) implies
Since the coefficient of each exponential in this expansion is bounded by The set of frequencies of is finite and contained in the interior of a closed pointed cone (contains no line) defined by the vertex of the Newton polytope Therefore the set of frequencies of the expansion satisfies
A combinatorial argument shows that there exists such that
Multiplication by the trigonometric polynomial preserves the asymptotic polynomial bound on the distribution of frequencies. This concludes the proof.
Lemma 18
There exist and such that and are tempered Radon measures and
Proof. Follows from preceding results by defining
where are the combinatorial coefficients.
Lemma 19
There exists a discrete subset and a sequence such that and are tempered Radon measures and
Proof. As and the corresponding coefficients converge. Therefore is a Cauchy sequence in the Hausdorff metric for every so converges pointwise to a discrete set and the coefficients converge to so that defined above satisfying the asserted properties.
Corollary 5
5.4 Proof of Theorem 3
Assume that is real-rooted and uniformly generic. If is unfolded then Proposition 10 implies that is a Fourier quasicrystal om If is folded then for every and choose a random unit vector and then for every define be the trigonometric polynomial whose frequency replaced by frequency and whose amplitude at equals the amplitude of at Define Since foldedness is defined by equalities, with probability one the trigonometric map is unfolded. Therefore there exist unit vectors such that is unfolded for every . Then Proposition 10 implies is a Fourier quasicrystal on Since and are uniformly bounded for sufficiently small the following analogue of Proposition 8 holds.
converges to uniform of bounded subsets of so the same argument used to prove Proposition 10 implies that is a Fourier quasicrystal.
6 Conclusions and Future Research
Section 2 provided a background in harmonic and functional analysis relevant to Fourier quasicrystals with a focus on almost periodicity. Proposition 2 implies that every Fourier quasicrystal on is a Bohr almost periodic measure. If this measure is supported on a discrete set and has coefficients in then it is a Bohr almost periodic multiset in If the spectrum of a Bohr almost periodic multiset in is finitely generated with rank then results in [44] show that its compactification is a union of continuous images of in the images are transverse to the foliation associated to the compactification and the homotopy of the associated maps is injective. Furthermore the spectrum can be computed by integrating certain differential -forms over these images. This provides a computational algorithm to determine if the multiset is a Fourier quasicrystal. All examples of multidimensional Fourier quasicrystals in Section 3 and elsewhere are of this toral type and the compactifications are intersections of with algebraic varieties of dimension in This raises two research questions:
Question 1
Is the spectrum of every Fourier quasicrystal on with coefficients in finitely generated?
Question 2
If the spectrum of a Fourier quasicrystal on with coefficients in is finitely generated with rank is its compactification in the intersection of with an algebraic variety of dimension in The result by Olevskii and Ulanovskii [66] imply that this holds for
If the answer to this question is yes then results in [44] give formulae for computing the density of the multiset which equals the density of real zeros of the associated trigonometric polynomial map where is determined by the variety and is determined by the compactification of the multiset. Proposition 4 implies that is real-rooted iff the density of real roots equals the density of all roots of The later is given by Bernshtein’s formula (59) since transversality implies that is generic. Section 4 provides a method to represent real-rooted uniformly generic trigonometric maps in the form where is determined by and is -stable. For Alon and coauthors [1] proved that can be chosen to be a Lee-Yang polynomial. Ruelle [74] completely characterized these polynomials. This raises the following question:
Question 3
Given a rank matrix with how can all -stable Laurent maps be characterized?
Acknowledgments The authors thank Sergey Favorov, Alexander Olevskii, Yves Meyer, and Peter Sarnak for sharing their knowledge about Fourier quasicrystals and Lang’s conjecture. We thank the reviewer for suggesting major revisions that improved accuracy and readability of the paper.