### v3

# Fourier Quasicrystals on

Abstract
This paper has three aims. First, for we construct a family of real-rooted trigonometric polynomial maps whose divisors are Fourier Quasicrystals (FQ). For these divisors include the first nontrivial FQ with positive integer coefficients constructed by Kurasov and Sarnak [47], and for they overlap with Meyer’s curved model sets [65] and two-dimensional [66] and multidimensional [67] crystalline measures. We prove that the divisors are FQ by directly computing their Fourier transforms using a formula derived in [50]. Second, we extend the relationship between real-rootedness and amoebas, derived for by Alon, Cohen and Vinzant [1], to the case The extension uses results in [10] about homology of complements of amoebas of algebraic sets of codimension Third, we prove that the divisors of all uniformly generic real-rooted are FQ. The proof uses the formula relating Grothendieck residues and Newton polytopes derived by Gelfond and Khovanskii [34] . Finally, we note that Olevskii and Ulanovskii [72] have proved that all FQ with positive integer weights are divisors of real-rooted trigonometric polynomials for but that the situation for remains unsolved.
2020 Mathematics Classification 52C23; 32A60; 32A27 111This work is supported by the Krasnoyarsk Mathematical Center and financed by the Ministry of Science and Higher Education of the Russian Federation (Agreement No. 075-02-2024-1429).
Contents
- 1 Introduction
2 Preliminaries
- 2.1 Notation and Conventions
- 2.2 Tempered Distributions and Radon Measures
- 2.3 Crystalline Measures and Fourier Quasicrystals
- 2.4 Bohr and Besicovitch Almost Periodicity
- 2.5 Bohr and Besicovitch Almost Periodic Multisets
- 2.6 Holomorphic Maps and Multiplicity of their Roots
- 2.7 Representation of Trigonometric Maps
- 2.8 Minkowski Geometry and Mixed Volumes
- 2.9 Genericity for Laurent Maps
- 2.10 Genericity and Uniform Genericity for Trigonometric Maps
- 2.11 Pontryagin Duality for Locally Compact Abelian Groups
- 2.12 Compactification of Groups, Functions, Measures. Multisets
- 3 Examples
- 4 Amoebas and Real-Rootedness
- 5 Sufficient Conditions for Fourier Quasicrystals on
- 6 Conclusions and Research Questions
1 Introduction
In 1974 Roger Penrose [75], [37] introduced his first aperiodic planar tilings with fivefold rotational symmetry.
In 1981 de Bruijn [12] described a ”cut and project method” to construct Penrose tilings by projecting a subset of a cubic lattice in five dimensions onto a plane.
In 1983 Paul Steinhardt and his student Dov Levine, inspired by these tilings, proposed a general concept of quasicrystals in a patent disclosure and published it in 1984 [54].
Working independently, in 1984 Dan Schechtman and colleagues [85] synthesized substances exhibiting Bragg diffraction peaks like crystals but whose icosahedral
symmetries were theoretically impossible for crystals. Schechtman was awarded the Chemistry Nobel Prize in 2011 for discovering these quasicrystals. In 2019 Steinhardt [87] described the international expedition he led to Kamchatka in 2011 that found natural quasicrystals in meteorites.
The cut and project method to construct models for both aperiodic tilings and quasicrystals are similar to the model sets studied a decade earlier by Meyer [60]. Also see [61, 62, 68]. The model sets are discrete but define Radon measures whose
spectrums, defined by their Fourier transforms, are supported on dense subsets.
Lagarias [48] proved that model sets are
Besicovitch but not Bohr almost periodic.
More recently Meyer
[63, 65, 66, 67] and Lev and Olevskii [55, 56, 57, 58] constructed crystalline measures - those whose supports and spectrums are both discrete, and a subclass called Fourier quasicrystals.
Favorov and colleagues
[15, 16, 17, 21, 24, 25, 26, 44] studied almost periodic sets and their relation to divisors of holomorphic almost periodic functions in tube domains.
In 2020 Kurasov and Sarnak [47] proved that the divisor of every real-rooted univariate trigonometric polynomial is a Fourier quasicrystal and constructed such polynomials from Lee-Yang polynomials. Shortly after in 2020 Olevskii and Ulanovskii [72] proved that every Fourier quasicrystal on is the divisor of a real-rooted univariate trigonometric polynomial.
In 2023 Alon, Cohen and Vinzant [1] proved that every real-rooted exponential polynomial (called trigonometric polynomial in this paper) is the restriction of a Lee-Yang polynomial. These three results in combination completely characterize Fourier quasicrystals with positive integer weights on
As summarized in the abstract this paper has three aims. Achieving these aims required using extensive notation and background knowledge. For this purpose we have included a long section titled Preliminaries. A key result is Fourier quasicrystals with positive integer coefficients are Bohr almost periodic multisets but Section 3.1 gives examples showing that the converse fails. For these examples the compactifications of multisets are not algebraic varieties. This fact supports our conjecture that the result of Olevskii and Ulanovskii referenced in the abstract holds for Fourier quasicrystals with positive integer weights on for More specifically, the support of the Fourier quasicrystals are the set of commin zeroes of a set of trigonometric polynomials of on where All examples describe by Alon, Kummer, Kurasov and Vinzant in [2] have weight and one example has In Example 3 in Subsection 3.2, whenever and all weights However, in Theorem 3 in Section 5, and the weights equal the multiplicities of zeros of trigonometric polynomial maps
We strove to make the paper self contained. However, a detailed explanation of the material in Subsection 5.2 would increase the length of this paper to over 100 pages so we merely referenced the seminal paper of Gelfond and Khovanskii [35].
2 Preliminaries
2.1 Notation and Conventions
Throughout this paper we use the following notation and facts.
means ’is defined as’, iff if and only if,
empty set, composition of maps, homeomorphism, isomorphism.
For sets number of elements in
For a locally compact Hausdorff () topological space
locally convex topological vector space of continuous complex-valued functions on with topology of uniform convergence on compact subsets . This is a Fréchet space iff is countably compact.
Banach space of bounded continuous functions with norm
its Banach subspace of functions converging to at infinity (for every there exists a compact such that on ).
its dense subspace of compactly supported functions,
space of uniformly continuous where is a uniform space (metric space or topological vector space).
space of complex-valued Borel (finite) measures on
space of complex-valued Radon measures on
are integer, nonnegative integer, positive integer, rational, real, positive real, and complex numbers.
are the real part, imaginary part, modulus of
Tuples are represented as column vectors.
transpose.
For maximal number of -independent elements in
-module generated by
if then
Lebesgue Banach space with norm for
open ball with center radius in
open ball with center radius in
Thus is congruent isometric to
is the Lebesgue measure of
set of by real matrices defining maps
maximum number of -independent rows of
is the general linear group over
is the identity matrix.
and are topological groups under multiplication.
and are compact subgroups.
is the identity and multiplication is denoted by juxtaposition.
is the homomorphism with kernel defined by
defines a homomorphism defined by
Thus defines the homomorphism by
Furthemore
algebra of Laurent polynomials consisting of sums
where the spectrum is finite
and is the amplitude corresponding to frequency
convex hull of is the Newton polytope of
set of vertices of
is the set of roots of
set of Laurent maps
is the set of roots of
induces a homomorphism
defined by
thus
algebra of trigonometric polynomials consisting of sums
where the spectrum is finite
and is the amplitude corresponding to frequency
convex hull of is the Newton polytope of
set of vertices of
is the set of roots of
set of trigonometric maps
is the set of roots of
is real-rooted if
If is real-rooted and is discrete then
is the divisor of Here is the point measure at defined in 2.2 and
is the multiplicity of defined in 2.6.
2.2 Tempered Distributions and Radon Measures
For a topological vector space
its dual space consisting of continuous linear functionals
space of Schwartz [83] functions all of whose derivatives exist and decay fast. Its topology is defined by seminorms
space of tempered distributions. Identify by
is the Fourier transform defined by
It is a continuous bijection and
For Fubini’s theorem ([79], p. 164) gives
This allows us to consistently define the bijection by
Let for
is a dense subspace of The
Fourier-Laplace transform of defined by
is the unique analytic continuation of to an entire function.
Lemma 1
implies there exists depending only on the support of and for every there exists such that
Conversely, (13) implies that
Proof. This is the Paley–Wiener-Schwartz theorem ([39], Theorem 7.3.1).
If then its total variation defined in ([79], p. 116),
is a positive Borel measure, and is a Banach space with norm
([79], Theorem 6.2).
The Riesz representation theorem ([79], Theorem 6.19) implies
that
The point measure at
for every Borel Therefore
where has the inductive limit topology ([88], p.515). A Radon measure is discrete if it
equals
where is discrete (has no limit points)
and are called its coefficients or weights. Lebesgue measure
on is a Radon measure but not a Borel measure.
The following is from ([4], Definition 2.1).
Definition 1
is a tempered Radon measure if
there exists
such that
Since is dense in tempered Radon measures are tempered distributions. Hörmander ([39], Theorem 2.1.7) proved that compactly supported positive distributions are positive Borel measures, hence every positive tempered distribution is a positive tempered Radon measure.
Lemma 2
For positive the following conditions are equivalent.
1.
is a tempered Radon measure.
2.
There exist such that
3.
There exists such that
Favorov ([19], Lemma 1) called property 2 slowly increasing and noted its equivalence to property 1. The equivalence of properties 2 and 3 follows from
Remark 1
If is Lebesgue measure then the Fourier inversion formula implies
hence The Dirac comb is clearly a Radon measure.
Lemma 3
2.3 Crystalline Measures and Fourier Quasicrystals
Definition 2
A Crystalline Measure (CM) on is a discrete tempered Radon measure whose Fourier transform is a discrete Radon measure. A Fourier Quasicrystal (FQ) on is a CM on such that both and are tempered.
Remark 2
Favorov [20] constructed a CM on that is not a FQ.
Lemma 3 implies that the Dirac comb is a FQ. For and the affine transformation on gives am affine transformation on discrete Radon measures If is a CM with then a computation gives
so is a CM. Furthermore if is a FQ then is a FQ. Clearly finite linear combinations of CMs (resp. FQs) are CMs (resp. FQs). Furthermore, if is a CM (resp. FQ) on and is a CM (resp. FQ) on then their product is a CM (resp. FQ) on
Definition 3
A CM on is called trivial if it is a finite linear combination of affine transformations of Dirac combs multiplied by trigonometric polynomials. Then its support is a finite union of arithmetic sequences and the CM is a FQ. A CM on for is called trivial if it is a finite union of affine transformations of multidimensional Dirac combs multiplied by trigonometric polynomials; or a finite linear combination of of affine transformation of products of (not necessarily trivial) CMs on lower dimensional spaces. A FQ is called trivial if it is a trivial CM. It is called nontrivial if it is not trivial.
2.4 Bohr and Besicovitch Almost Periodicity
The theory of uniform almost periodic functions on was pioneered in three papers [7] by Harald Bohr and extended to mean-squared almost periodic functions in [6] by Abram Besicovitch. We summarize their results extended to define almost periodic measures and relate them to tempered distributions and Fourier quasicrystals in Propositions 1 and 2.
Definition 4
is relatively dense if there exists such that is almost periodic ([8], p. 32, Main Definition) if for every the set
is relatively dense. is called the set of -almost periods of
We call these Bohr almost periodic functions, denote the set of them by and record that
1.
([8], Theorem I)
2.
([8],Theorem II)
3.
([8],Theorem III) is closed under addition, so
4.
([8],Theorem IV) is closed under multiplication
5.
([8],Theorem V) is closed under uniform closure.
6.
([8], p. 39, 44) every has a mean value
(17) where the value is indepedent of and convergence is uniform with respect to
- 7.
8.
([8], p. 60-65) This uniqueness is equivalent to Parseval’s equation
(19) -
9.
([8], p. 80-88, Fundamental Theorem) uniform closure of
(19) implies that the spectrum
is countable. For we call the set of translates of its orbit and the uniform closure its orbit closure. Bochner [9] proved
Besicovitch [6] defined generalized almost periodic functions completion of with respect to the norm
Clearly Lagarias [48] proved that functions associated with quasicrystals and Meyer’s model sets are in Meyer [62] explains this result iin detail.
Remark 3
A subset of an abelian topological group is syndetic if there exists a compact such that For syndetic means relatively dense. These concepts play a central role in topological dynamics where thier relation to minimal systems was pioneered by Gottschalk and Hedlund ([36], Chapter 4) . For and define the translation of by and the orbit of , Then is a uniformly equicontinuous group of transformations of ([36], 11.08) so for every and every extends to a map ([36], 4.44) implies that iff is compact. This gives Bochner’s [9] characterization of almost periodicity: is Bohr almost periodic if every sequence of translations of has a subsequence that converges uniformly. Moreover ([36], 4.48) implies that the following conditions are equivalent:
1.
2.
is a compact minimal closed subset of invariant under
3.
there is a unique group structure on which makes it a compact topological group and the map defined by has a dense image.
If define and by for Then The Stone-Weierstrass theorem ([78], A14) implies that the algebra of trigonometric polynomials on is dense in so can be uniformly approximated by hence can be uniformly approximated by which implies Bohr’s fundamental theorem. These results illustrate the utility of topological dynamics
If and
then
where is defined in (5). The right side of (24) converges uniformly independent of the ordering of terms.
([8], p. 89, Convergence Theorem) shows that this always occurs whenever
is linearly independent over
If (23) fails then
the right side of (24) must be computed as a limit of Fejer sums as explained in ([8], p. 66).
The convolution of and
is in and
Definition 5
A Radon measure is a Bohr (resp. Besicovitch) almost periodic measure if for every the convolution is a Bohr (resp. Besicovitch) almost periodic function. A Bohr (resp. Besicovitch) almost periodic distribution is a tempered Radon measure such that for every is a Bohr (resp. Besicovitch) almost periodic function.
If is a Bohr almost periodic measure define by
where and (25) ensures that is well defined. The spectrum
is countable since there exist such that for all Meyer ([63], 2nd sentence) calls Radon measures that are are countable linear combinations point measures purely atomic measures. Meyer [67] and Favorov [22] define a tempered distribution a to be a Poisson measure if both and are purely atomic measures.
Lemma 4
If is a Bohr almost periodic measure then iff
Then
is a purely atomic measure. Therefore, if is a Bohr almost periodic purely atomic measure, then is a Poisson measure iff (28) holds.
Proof. Meyer proved this ([64], Theorem 5.5).
Remark 4
Favorov ([18], Lemma 1) proved that if is a tempered measure, is a pure point measure, and for some then is an almost periodic distribution and for every converges absolutely.
We record the following observation without proof.
Remark 5
If is a Bohr almost periodic measure on and (28) fails, then there exists such that and is differentiable and has derivative except on where has a jump equal to at every In higher dimensions is more complicated.
Definition 6
A Radon measure on is translation bounded if there exists such that
Since is contained in a union of balls of the form Lemma 2 implies that every translation bounded Radon measure is a tempered Radon measure.
Proposition 1
If then is a Bohr almost periodic measure iff is translation bounded and is a Bohr almost periodic distribution.
Proof. Let be a Bohr almost periodic measure. Let be the Banach space of consisting of whose supports are contained in and let For every define by
Clearly each is a bounded lineear functional on since
Furthermore, since
Therefore the Banach-Steinhaus theorem ([79], 5.8) implies that there exists such that
hence is translation bounded. Let We conclude the proof of the only if part by proving that Let be a sequence such that equals on and for Then and converges uniformly to Since each and is closed under uniform limits,
To prove the if part assume that is a translation bounded measure and a Bohr almost periodic distribution. Let We conclude the if part of the proof by showing that
Let such that Then is a bounded continuous function and is the uniform limit of hence
Sergey Favorov proved ([23], Theorem 1) that every non-negative Poisson measure is a Bohr almost periodic measure. Proposition 1 implies that his result gives the following result as a special case. We provide a proof to make the paper more self-contained.
is
Proposition 2
Every positive FQ on is translation bounded and is a Bohr almost periodic distribution hence is a Bohr almost periodic measure.
Proof. Let be a positive FQ so where is discrete, and is a tempered distribution. Therefore there exists such that
Let Then where hence
Since there exists such that for Therefore there exist with
This proves and has an absolutely converging Fourier-Bohr series so is a Bohr almost periodic distribution. Choose values in whose support is in and equals on Since it is bounded and positive implies hence
so is translation bounded. Proposition 1 then implies that is a Bohr almost periodic measure.
2.5 Bohr and Besicovitch Almost Periodic Multisets
Krein and Levin [45] introduced almost periodic multisets in to study zeros of holomorphic almost periodic functions. Favorov and Kolbasina [24] studied them in In this subsection we discuss them in to study zeros of holomorphic maps in Subsection 2.7.
Definition 7
A multiset in is a pair where is discrete and has support We let denote the set of multisets in and denote its subset where on If we say is a multiset in
can be represented by a sequence where for every for values of Two sequences and represent the same multiset iff there exists a permutation
such that
This defines equivalence classes of –valued sequences.
Define the distance between multisets
where is represented by a sequence is represented by a sequence and ranges over all permutations of We observe that the metric space is complete since every Cauchy sequence converges.
Remark 6
The restriction of to is the Hausdorff distance explained in ([69], p. 279, Exercise 7)
For define the translation by
Since translations are isometries, the dynamical system is uniformly equicontinuous. Following the logic used to analyze Bohr almost periodic functions in Subsection 2.4, we define to be Bohr almost periodic if for every the set of -almost periods
is syndetic. Therefore is Bohr almost periodic iff its orbit closure is compact and then admits a unique compact group structure such that the map defined by is a continuous homomorphism with a dense image. Thus the pair is a compactification of in Definition 15. Define the sum of multisets by
Associate to a multiplicity set the unique measure
Kolbasina and Favorov [24, 44] proved that a multiset in is Bohr almost periodic iff is a Bohr almost periodic measure and then its density
exists, is independent of and the sum converges uniform in We define a multiset in to be Besicovitch almost periodic if is a Besicovitch almost periodic measure. Favorov[17] characterized Besicovitch almost periodic multisets in by their translation properties.
Definition 8
For define
is uniformly discrete if and is relatively uniformly discrete if it is the sum of a finite number of uniformly discrete multisets. The index of is
We observe that is relatively uniformly discrete iff and then equals the sum of uniformly discrete multisets. Furthermore, all Bohr almost periodic multisets in are relatively uniformly discrete.
Remark 7
Assume that is uniformly discrete, Bohr almost periodic, and and let be the associated compactification of Then ([50], Theorem 5) implies that and called the compactification of is homeomorphic to a finite union of pairwise disjoint submanifolds of each homeomorphic to and embdede transverse to the foliation induced by and inducing an injective homotopy.
2.6 Holomorphic Maps and Multiplicity of their Roots
A domain is a nonempty open connected set. A map is holomorphic if each component is holomorphic. A root of is isolated if there exists such that and Assume that is an isolated root of For and the multiplicity of equals its logarithmic integral
where is the one-cycle consisting of a circle with center at bounding a closed disc contained in and containing only the root and oriented so is positive (counterclockwise orientation). For the multiplicity can be computed by a Bochner-Martinelli integral over a sufficiently small -dimensional sphere centered at ([89], p. 16, Eqn. 1), ([84], p. 286, Eqn. 3). Alternatively, its multiplicity can be computed by an integral
over the -cycle
where is a neighborhood of with are sufficiently small to ensure that and is oriented so that the differential form We call a Grothendieck cycle since (44) is the Grothendieck residue of the Jacobian ([89], p. 16). The Bochner-Martinelli integral representation of multiplicity gives:
Lemma 5
(Rouche) If is a bounded domain with Jordan boundary and if are continuous, holomorphic on and on then has as many zeros in as
Proof. Follows from ([84], Theorem 2, p. 288).
Definition 9
A (horizontal) tube domain is a subset where is nonempty, open, connected and bounded.
The set of bounded holomorphic functions on is a Banach space with norm For discrete define the multiset where and is supported on and for We also identifty with the Radon measure
Definition 10
A holomorphic map is Bohr almost periodic if it is bounded and for every there exists a syndetic such that
denotes the Banach space of holomorphic Bohr almost periodic maps
Roots are counted with multiplicities. If and is discrete then is a multiset and a Radon measure and we define the lower density of by
Corollary 1
Proof. Let with
Since
there exists a syndetic
with
has roots in
and
so Lemma 5 implies that has roots in
and has roots in
Since
is syndetic it must contain a subset with lower density
such that the balls are pairwise disjoint. This implies that
and concludes the proof.
For define the compact set and the closed set
Corollary 2
If and is discrete for every then for every
Proof. Assume to the contrary that there exists with Then there exists a sequence with Since both and are compact, we may replace with a subsequence satisfying the following property: there exist and such that
Therefore
Since is discrete there exists such that
Lemma 5 implies that for sufficiently large has a root Since and
We may also choose so that hence since
This contradicts (48) and concludes the proof.
Corollary 3
If and for every the set is discrete, then
1.
If is compact, then
2.
is a Bohr almost periodic multiset.
3.
is contained in the group generated by
Proof. To prove 1. assume to the contrary that there exists a sequence such that
Since
is compact so can be replaced by a
subequence such that
Since is bounded the partial derivatives of its components
are bounded hence has an infinite number of zeros in
This contradicst the assumption that is discrete and
concludes the proof.
To prove 2. we first observe that 1. implies there exists
and such that each connected component of
is the union of at most sets
Let be interiors of these connected components. The diameter of each is Corollary 2 gives
There exists a syndetic with
Therefore Lemma 5 implies that
Since the diameter of each is this implies
hence is a Bohr almost periodic multiset.
To prove 3. define and observe that 2. implies that
the pair where is defined by is a compactification of as is
Definition 15. Also the pair
where is defined by is a compactification of Lemma 10 implies that the spectrum
generates the spectrum of the comactification and hence is contained in it. A similar argument gives that generates the spectrum of the compactification and hence is contained in it.
Lemma 5 implies that the homomorphism defined by
is uniformly continuous with respect to the metric
on and the metrix on
Therefore extends to a
continuous homomorphism
and
is an equivariant map between minimal dynamical systems hence
and
Therefore
We note that the proof of Lemma 10 does not require any result from Corollary 3 so the proof of 3. does not employ circular reasoning. The following example shows that the inclusion in 3. can be strict. If is a tube domain containing then and Poisson’s summation formula implies
2.7 Representation of Trigonometric Maps
For we write if the exists such that This is an equivalence relation.
Definition 11
is minimal if
Proposition 3
If and then
where is defined in (1), and
If and then there exists a minimal and such that For every minimal and that satisfy there exists such that and
Proof. The first assertion follows from a direct computation. Let and whose columns form a -basis for and For every there exists a unique such that Then for every entry of (6) implies that
where is defined by
and is defined following (2). Therefore where Since Since the column of are -independent, so is minimal. To prove the third assertion note that hence there exist with and and Then Since the columns of are -independent, hence hence Let Then (3) implies
Lemma 1 implies that has a dense image hence so
Corollary 4
corollary4 If and is a tube domain then
Proof. For and hence
Proposition 3 implies there exists and such that Let is a component of and the corresponding component of Then Let For the triangle inequality gives
where
(51) gives
acts on to give a dynamical system
This action is uniformly equicontinuous and since is compact every orbit is almost periodic. Theregore for every there exists a syndetic such that
For every choose
(52), (54) and the fact that implies gives
2.8 Minkowski Geometry and Mixed Volumes
For any compact convex –dimensional volume of The Minkowski sum of compact convex and is In [5] Bernshtein defined the mixed volume of an -tuple of compact convex sets
and noted that it differs from the classical definition by the factor We will use Bernshtein’s definition in the remainder of this paper.
Lemma 6
iff there exists points such that is linearly independent. Then
Proof. ([82], Theorem 1.5.7, Theorem 5.1.8).
Lemma 7
If the vertices of are in then
Proof. ([14], Theorem 3.9).
2.9 Genericity for Laurent Maps
The fundamental theorem of algebra ensures that if then
has roots in counting multiplicity. The situation for Laurent maps is more complicated even for a system of two linear equations in two variables. Consider where
Then where is the triangle with vertices
A computation shows iff
and then is called generic.
Otherwise either or is unbounded and has no isolated points. We note that each of the three inequalities that characterize genericity involves amplitudes associated with vertices on one side of
For is analytic so is finite iff it is compact
([11], Proposition 1, p. 31) and genericity of was
defined by Bernshtein [5] who proved
Lemma 8
If is generic then is finite and
If is not generic then either or the set is noncompact and its subset of isolated roots satisfies
and
Fix and for let be a polytope whose set of vertices The set of with is parameterized by
where and
We define genericity for Laurent maps.
For and
define
and for
Definition 12
is generic if for every
is a face of that is orthogonal to and has dimension Therefore there are only a finite number of Furthermore, Bernshtein ([5], p. 184) noted that each
”depends essentially on smaller number of variables” and therefore
”does not have any zeros in the general case.” He showed that the set of non generic are parameterized by a proper algebraic subset of the set of parameters described above.
Therefore if is generic and satisfies
and the amplitudes (coefficients) of are sufficiently close to the amplitudes of then is generic.
We illustrate genericity with an example more complex than the previous example. Consider the set of
where
and
is the square with vertices These set of Laurent polynomials are parametrized by
Then
where all coefficients are in
If and then
is generic when all Clearly the last four Laurent maps above have no roots so a calculation shows that fails to be generic on the algebraic subset of that is the union of four hypersurfaces defined by
If is generic then calculation shows that has two roots and the mixed volume
2.10 Genericity and Uniform Genericity for Trigonometric Maps
Every nonzero has the form
where and
are real. Then for sufficiently large
dominates the other terms in and for
sufficiently large dominates the other terms in In both cases is bounded. Moreover, is uniformly bounded for sufficiently close to and sufficiently close to Since is a nonzero univariate holomorphic function, iff otherwise it is an infinite discrete set. Corollary 4 implies is Bohr almost periodic. Hence corollary 3
implies is Bohr almost periodic so (74) implies that
its density exists. A contour integration argument gives
The situation for trigonometric maps is much more complicated and we must impose stringent conditions on
to ensure that is bounded, is discrete and its density exists, and that these properties are preserved under sufficiently small perturbations.
We define genericity for trigonometric maps.
For and
define
and for define
Definition 13
is generic if for every
Let and for define
We observe that if then We note that and only the arguments but not the moduli of the amplitudes of differ from those of
Definition 14
is uniformly generic if is generic for all
Gelfond [33] announced that if is uniformly generic then is discrete, is bounded, and there exists with
(63) implies that the density
Remark 9
Gelfond deals with holomorphic Bohr almost periodic maps on vertical tube domains and counts roots whose imaginary parts lie in a cube He calls his condition properly unfolded. We avoid this term here because we reserve it for a condition that for every consists of a single vertex for some This condition implies uniform genericity but is stronger. Gelfond’s result applies to trigonometric maps since corollary 3 implies they are Bohr almost periodic. For trigonometric maps the indicator diagrams of the components are their Newton polytopes so Gelfond’s condition ([32], System 3, p. 155) applies and the requirement that be uniformly generic is stated in ([32], (c), p. 156) ”it is required that this closure does not contain consistent systems” meaning and Gelfond’s paper does not provide a proof but references his preprint [31]. He also references a paper of Hovanskii [41], in which 4 out of 7 references are preprints, and a paper by Kazarnovskii [42] in Russian. Kazarnovskii’s paper [43] derived asymptotic distribution of roots for exponential sums whose exponents are not necessarily purely imaginary and thus not trigonometric polynomials.
Proposition 4
If is uniformly generic then there exists a tube domain such that for all sufficiently close to and is a Bohr almost periodic multiset contained in Moreover, is real-rooted iff
Proof. Follows from Corollaries 1, 3, and 4.
The following result extends Levin’s result ([53], p. 268, Lemma 1) for
Proposition 5
If is uniformly generic and is a tube domain containing is the set of such that whenever or Then
Proof. Follows from Corollaries 2 and 4.
2.11 Pontryagin Duality for Locally Compact Abelian Groups
Let be a locally compact abelian group. A continuous homomorphism is called a character of The Pontryagin dual of is the set of characters of equipped with pointwise multiplication and the topology of uniform convergence on compact subsets. is a locally compact abelian group with identity the constant function with value We observe that
is discrete (resp. compact) iff is compact (resp. discrete). If is the additive group of a field then is isomorphic to though not in a unique way. is connected iff is torsion free (no nonidentity elements has finite order). If and are locally compact topological groups and is a continuous homomorphism then its Pontryagin dual defined by
is a continuous homomorphism. Moreover, is injective (resp. has dense image) iff has dense image (resp. is injective). Pontryagin’s duality theorem ([78], Theorem 1.7.2) implies that the inclusion defined by
is a isomorphism of topological groups, so we identify with and if we identity with Therefore characters on separate points of and iff If is a closed subgroup of a locally compact abelian group define
Pontryagin duality gives
If is a closed subgroup of then the quotient is a locally compact abelian group and the map
defined by is a continuous surjective homomorphism with kernel
Hence iff iff there exists iff with
Therefore if is a closed subgroup of then iff there exists
such that and if
is a closed subgroup of then iff there exists
such that
The following closure results are used throughhout the sequel of this paper.
Lemma 9
Assume and is defined by Then
1.
2.
3.
iff
4.
iff
Proof. 1. follows from (6) since iff there exists with
and the last inclusion holds for some iff
2. follows since kernel implies
so
iff
3. follows since iiff there exists with
and the last equality holds for some iff
4. follows since
iff there exists with
iff iff
2.12 Compactification of Groups, Functions, Measures. Multisets
Definition 15
A compactification of a locally compact abelian group is a pair where is a compact group and is a continuous homomorphism with a dense image. Clearly is abelian and if is connected then is connected. The Bohr compactification is when where equals with the discrete topology and and is the identity homomorphism.
The Bohr compactification has the following universal property: If is any compactification of then there exists a surjective homomorphism such that If is a compactification of we define its spectrum If is the group generated by with the discrete topology then Pontryagin duality gives Let the algebra of finite linear combinations of characters on Since separates points and is closed under complex conjugation, the Stone-Weierstrass theorem ([78], A14) implies that it is dense in For let the discrete subgroup of generated by
Lemma 10
If and then and where integration is with respect to Haar measure on Let be the inclusion map. If then where and is a compactification of Furthermore and there exists a unique such that
Proof. Since is dense in can be uniformly approximated hence can be approximated by so The integral identity occurs because if else Since is discrete is compact and since is injective has a dense image hence is a compactification of Pontryagin duality implies that and hence Then so there exist a sequence with and Since there exists a sequence such that Since has dense image Since converges it is a Cauchy sequence hence is a Cauchy sequence so converges to a function Then
Definition 16
Let denote the set of Borel measures on a compact abelian group The Riesz Representation theorem ([78], Theorem 6.19) implies that is the space of continuous linear functionals on For we define its Fourier transform by
and its spectrum
Definition 17
An approximation of the identity on is a sequence that is nonnegative, and for every open set containing there exists such that support Then converges weakly to
Proposition 6
If is a Bohr almost periodic measure on then there exists a compactification of and such that
The measure is called the trace of with respect to and it is unique.
Proof. Let be the discrete subgroup of generated by then construct the compactification of as in the proof of Lemma 10. Let be an approximation of the identity on Since is translation bounded
Since is a Bohr almost periodic measure there exists a sequence of such that Furthermore are nonnegative and so Therefore the Banach-Alaoglu Theorem ([80], 3.15) implies that there exists a subsequence of that converges weakly to a measure on Replacing the sequences with this subsequence and applying (25) and Lemma 10 gives
Since is dense in is uniquely determined
Remark 10
The trace of with respect to the Bohr compactification equals the image of under the Bohr map ([3], Section 7).
3 Examples
Example 1 constructs a trivial Fourier quasicrystal that is not uniformly discrete, hence not Delone. Example 2 constructs two model sets that are Besicovitch but not Bohr almost periodic. Moreover, neither of their associated measures are Poisson because their Fourier-Bohr transforms do not satisfy (28). However, their sum is a Bohr almost periodic Poisson measure. Example 3 constructs a family of uniformly discrete Bohr almost periodic subsets whose associated measures are Fourier quasicrystals. The algebraic varieties they use are generally not complete intersections (codimension and defined by the common zeros of Laurent polynomials). In this respect they differ from the construction of Fourier quasicrystals in Section 5 but are similar to the construction by Alon, Kummer, Kurasov and Vinzant [2].
3.1 Almost Periodic Multisets
Example 1
We record the following observations without proof: these are FQs on that by Definition 3 are trivial, their indices defined by (42) satisfy Since these multisets are Bohr almost periodic, their orbit closures are compact groups. Furthermore, computations give
Example 2
Parametrize by with coordinates Let and define the homomorphism by
Since is irrational, has a dense image so is a compactification of Let and define three -chains in
Since each is transverse to the foliation on whose leaves are the cosets of Therefore each is a uniformly discrete subset of so we can define the multiset and associated pure point Radon measure by
Clearly and
and are Meyer model sets [68], ([50], Definition 2). Therefore, as noted in Subsection 2.4 they are Besicovitch almost periodic but not Bohr almost periodic. However, since is a cycle (closed loop) is Bohr almost periodic. We will prove that for every with
Lemma 4 and (67) imply that and are not Radon measures and is a Radon measure. Therefore and are not Poisson measures and is a Poisson measure.
Our proof combines results in [50] about Bohr almost periodic sets and results in [46] about uniform distribution.
([50], Remark 4) implies that
and for
hence
Since the quantities in (69) and (70) never vanish,
which is dense in by Lemma 9.
For
where hence
Since
the second term in the right side of (73) equals hence (67) holds for
A calculation shows that the first term in the right side of (73) never vanishes hence
is dense in
Define by
Since Lemma 9 implies and ([46], Example 6.7) shows that the sequence is uniformly distributed. This means that for every nonempty open
where Haar measure of Since is non empty and open, hence the density of the sequence is positive. We observe that (74) implies
Choosing
gives iff there exists with
and
Therefore the second term on the right side of (71) is hence (67) holds for The proof for is similar.
3.2 Fourier Quasicrystals
We use results in [50] to construct Fourier quasicrystals on Let be a uniformly discrete Bohr almost periodic set of toral type and be its associated Bohr almost periodic measure. (The abstract of [50] says incorrectly that has toral type if the Fourier transform of equals zero outside of: a rank subgroup of It should say: a generated by elements subgroup of The group of dyadic rational numbers has rank but is not finitely generated.) The results in [50] imply:
1.
the associated compactification of equals where iff is a lattice subgroup of where
2.
where is an -dimensional submanifold of transverse to the foliation of whose leaves are the cosets of Each connected component of is homotopic to a subgroup of isomorpic to and the homotopy is transverse to the foliation.
3.
the Fourier transform is zero outside of and
(76) where is the unique differential -form transverse to the folliation and The density is determined by and the homotopy classes of the connected components of
These results imply that if has connected components, then is a union of irreducible Bohr almost periodic sets (meaning that they are not the union of Bohr alost periodic subsets). Furthermore, since (76) implies that and hence is a Fourier quasicrystall iff is bounded by a polynomial in
In this manuscript we use a different notation from the notation used in [50].
We denote the Fourier transform of a tempered distribution
by and denote the Fourier-Bohr transform of an almost periodic measure by
[50] denoted the Fourier transform of a tempered distribution
by and denoted the Fourier-Bohr transform of an almost periodic measure by
We define the Fourier-Bohr transform of an almost periodic set (or multiset) to be the Fourier-Bohr transform of its associated measure. Therefore, if is the measure corresponding to in the example below, then
in (83) equals
in the notation used in [50].
Example 3
Fix with distinct with and define
For every define by
and
The assumptions on definitions of and and Lemma 9 imply that is a compactification of and
Lemma 11
Every is injective. is a subgroup of transverse to the foliation on is a lattice subgroup of whose density is determined by and The family of maps is a homotopy so these maps induce identical maps on fundamental groups:
namely the map defined by There exists such that for every is transverse to the foliation on (whose leaves are the cosets of is a uniformly discrete Bohr almost periodic set that satisfies and is a Bohr almost periodic perturbation of hence its density The spectrum is irreducible. is nontrivial as defined by Definition 3.
Proof. implies there exists with Furthermore (78) implies
Therefore
hence determines so it
is injective. Therefore is a subgroup of
is an isomorphism of onto and is transverse to the foliation on because
Since is continuously differentiable
function of there exists such that for every
is transverse to the foliation of
The remaining assertions follow from the theory
developed in [50]. In particular
is irreducible because is connected. is
nontrivial because if it was an affine transformation of the product of
Bohr almost peridic sets on lower dimensional Euclidean spaces, then would either be disconnected or have dimension
We observe that where is the subvariety defined by the set of common zeros of Laurent polynomials:
In general, if then is not defined by a subset consisting of of the equations in (82), hence is not a complete intersection. Furthermore,
This inclusion implies real-rootedness of related trigonometric maps.
A subset of is called an infinite arithmetic progression if it equals for and
Lemma 12
If then does not contain an infinite arithmetic progression.
Since are rationally independent is injective. Assume to the contrary that and contains an arithmetic progression Then is contained in a finitely generated and thereore finite rank subgroup of Define Then is a finite rank subgroup of Therefore the solution of Lang’s conjecture ([13], Theorem 10.10.1) implies is contained in the finite union of translates of torus subgroups of of dimension Since are distinct, does not contain the translate of any torus subgroup of dimension and hence is finite. Since is injective, is finite. This contradiction concludes the proof.
Remark 11
Lang [49] conjectured that if is a finite rank subgroup of and is a proper variety, then is contained in a finite number of cosets in of subgroups of isomorphic to where Liardet [59] proved this for and Laurent [51] proved this for Ritt [76] proved that if and equals on then there exists such that This result implies that the set of real zeros of is a union where does not contain an infinite arithmetic progression, and or is the union of a finite number of infinite arithmetic progressions.
Theorem 1
is a Fourier quasicrystal for every
Proof. The theory in [50] implies that the Fourier-Bohr transform of equals zero outside the spectrum and
where is the -cycle obtained by orienting by the counterclockwise orientation of and is the differential -form
Therefore and Since
it suffices to prove that there exists a polynomial such that
Since and are convergent power series in nonnegative powers of whenever all The fact that the Radon measure associated with is real-valued implies that
Therefore whenever all If then
Let If then either all or all and Then
satisfies inequality (86) and the proof is concluded.
Remark 12
For
is equivalent (under the automorphism ) to the Laurent polynomial that Kurasov and Sarnak used ([47], Equation 41) to construct the first nontrivial FQ on
4 Amoebas and Real-Rootedness
4.1 Background
In 1952 in their study of phase transitions in statistical physics Lee and Yang introduced a family of and proved ([52], Appendix II) that
so has all its zeros in the unit circle. This family is
where is symmetric with and is summed over all subsets of
Subsequently more general polynomials satisfying (91) were studied and named Lee-Yang polynomials.
In 2010 Ruelle completely characterized multiaffine Lee-Yang polynomials (degree in each variable) ([81], Theorem 3), higher degree in each variable Lee-Yang polynomials ([81], Section 10), and interpreted the latter as temperature dependent partition functions involving highr spins
([81], Section 11).
In 2005 Passare and Tsikh ([74], Theorem 2) interpreted (91) in terms of the amoeba concept to prove that polynomials defined by (92) satisfy (91). To explain this concept, define the logarithmic moment map
is a homomorphism from onto with kernel A subset of is called an algebraic variety if it is the set of common zeros of a set of Laurent polynomials. By the Noetherian property of the ring of Laurent polynomials, this set of polynomials can be chosen to be finite. is called a hypersurface if the set has one element and then has codimension is called a complete intersection if it equals the intersection of hypersurfaces and has codimension The amoeba of an algebraic variety is
Gelfand, Kapranov and Zelevinsky ([30], p. 194, Definition 1.3) defined amoebas for hypersurfaces and proved ([30], p. 195, Corollary 1.6) that the complement of the amoeba of a hypersurface is a finite union of open convex subsets. Several researchers [28, 29, 73] extended their work.. Clearly is a Lee-Yang polynomial iff or equivalently
In [1] Alon, Cohen and Vinzant used amoebas to prove that every real-rooted admits the representation
where
is a Lee-Yang polynomial, and
We observe that (95) holds if is replaced by a matric sufficiently close to a condition we call -stable.
This section proves Theorem 2, which represents uniformly generic real-rooted trigonometric maps by -stable Laurent maps. Here
This concept generalizes the concept of Lee-Yang polynomials.
4.2 -stable Laurent Maps
Definition 18
For a rank matrix is -stable if
for all in some open neighborhood of
Lemma 13
If have rank and is -stable, then for some stable Here the equivalence relation was defined in the first paragraph in Subsection 2.7. For we can choose and to be a Lee-Yang Laurent polynomial.
Proof. acts on the Grassmann manifold consisting of -dimensional subspaces of An element in is rational if it has the form where has rank The existence of the Smith normal form [70, 86] for integer matrices implies that acts transitively on the set of rational elements in Since the set of rational elements is dense, and the compact group acts transitively on every orbit in under the action of is dense. The hypothesis implies that there exists an open such that and
Since every orbit is dense, there exists such that so is open and Define Then
Since for all in some open neighborhood of is -stable. This proves the first assertion. To prove the second assertion assume that is nonzero and that a Laurent polynomial is -stable. Then there exist an open and such that and
The Smith normal form implies that there exists whose first column equals By adding sufficiently large positive integer multiples of to the last columns of we construct such that
Define Then for every
hence is a Lee-Yang Laurent polynomial.
4.3 Proof of Proposition 7
The following result is an extension to Laurent maps of the result for Laurent polynomials by Alon, Cohen and Vinzant ([1], Corollary 3.3).
Proposition 7
If and has rank and is uniformly generic, then
Proof. Define
Since is a variety and it suffices to prove that
If then hence
so
Since is closed,
Proving the reverse inclusion is more challenging. We use proof by contradiction. Assume that there exists
Since is closed there exists such that
therefore does not vanish on the set
Proposition 4 implies rank hence is bounded. Since is invariant under translation by elements in is a tube domain and Corollary 4 implies that Moreover the Hausdorff distance
Proposition 5 implies there exists such that
(106) and the fact that is surjective implies there exists such that
(106) implies there exists such that
and (112) and (113) implies therefore (111) implies
For define by
(111), (112), (113) and (115) imply
(104) and (114) imply there exists a sequence such that
Therefore
and the convergence is uniform over Choose sufficiently large that
so has no roots in This contradicts (116) and concludes the proof.
4.4 Proof of Theorem 2
Theorem 2
If is uniformly generic and real-rooted and then there exists a rank matrix with and an -stable with
Remark 13
The proof of Theorem 2 requires the following result.
Lemma 14
If and rank then there exists such that
for sufficiently close to
Proof. Let and let be the boundary of and The hypothesis implies that for every the open ray
Since each is a finite union of open convex subsets there must exist and such that the line segment is contained in one of these open convex subsets. Therefore there exists an open cone such that
and
Since is an open cover of there exists a finite collection whose union contains Define and Then
hence if is sufficiently close to then
The proof of Lemma 14 is concluded since
We complete the proof of Theorem 2 by contradiction. Assume that is uniformly generic and real-rooted and let be the representation in Proposition 3. Since is generic, rank Assume to the contrary that is not -stable. Then Lemma 14 implies that for every neigborhood of there exists and We may assume that is sufficiently small is uniformly generic so there exists with
Proposition 7 then implies
Let and be spheres radii and in oriented in a positive direction to define cycles in and in Then the homology classes in because the point lies between the spheres. However, these cycles are homotopic in as the following argument shows. Define
Then and define homotopies that deforms the -cycles and to homologous cycles in Therefore the cycles and are homologous in so in This means that the homology inclusion
is not injective. This violates ([10], Theorem 1) and concludes the proof.
5 Sufficient Conditions for Fourier Quasicrystals on
This section proves the following result:
Theorem 3
If is real-rooted and uniformly generic then is a Fourier Quasicrystal on
5.1 Approximation Strategy
Since is dense in it suffices to prove the existence of a discrete and such that both and are tempered Radon measures and
Assume that is real-rooted and generic and Theorem 2 implies there exists a rank matrix and an -stable Laurent map such that Define a sequence
where converges to Since is -stable we may assume that each is real-rooted and generic.
Lemma 15
Proof. Since each component of (resp. equals (resp. ) times the corresponding component of it follows that
where convergence is componentwise with respect to the Hausdorff metric. The conclusion follows (63).
Proposition 8
Proof. For let be the characteristic function of (11) implies
Let be a tube domain containing and are Bohr almost periodic in and converges to uniformly on compact subsets of Lemma 5 implies that converges Therefore the first term on the right above converges to implies Therefore
converges to as since has fast decay. Similarly the third term on the right above converges to
as
Express
where
is the least common multiple of the denominators of the entries in
and
(3) gives
where
Since and hence is generic Lemma 8 implies is finite. Then since is real-rooted so there exists a finite with
For define
If then and the multiplicities satisfy
since and multiplicity is preserved under the holomorphic local homeomorphisms. Therefore
The Poisson summation formula Lemma 3 and its consequent (15) imply
Definition 19
For and define by
Define differential -forms on
where and
Define -chains in
where is the Grothendieck cycle defined in (45) with replaced by
Lemma 16
5.2 Grothendieck Residues and Toric Geometry
In ([35], 1.2) Gelfond and Khovanskii introduced the following concept:
Definition 20
Let be an -tuple of polytopes in Then each face of their Minkowski sum is a Minkowski sum where each is a unique face of called a summand of The tuple is called unfolded if every face of has at least one vertex among it summands. Otherwise it is called folded.
The Laurent maps in in 2.7 have Newton polytope tuples where and are identical squares. Therefore is a square each of whose sides has summands consisting of sides not vertices. Therefore is folded. We record the following observation without proof. is unfolded iff for every there exists and a vertex such that
It is easy to show that if and the components of
satisfy (143) then is generic.
If a tuple of polytopes is unfolded then to each vertex of their sum is associated an integer called its combinatorial coefficient, see [33, 34, 35].
If is a trigonometric map and is unfolded, then for sufficiently close to (with respect to its frequencies and corresponding amplitudes) is unfolded
and whenever is a vertex of the sum of the polytopes in
and is the vertex of the sum of the polytopes in
that is closest to
We may assume that each is unfolded so is unfolded since its entries differ from those in by a constant multiple.
For every define the Newton cycle oriented so
is positive.
Proposition 9
If is unfolded and then for every there exists such that
1.
2.
In the restriction of to the monomial corresponding to dominates the other monomials in so that for some
(144) such that for all
3.
is homologous to the Newton -chain defined by
(145) where is the combinatorial coefficient for the vertex
4.
Then Lemma (16) and Stokes theorem imply
Proof. These results are derived in ([35].
5.3 Proof of Proposition 10
Proposition 10
If is real-rooted and is unfolded, then is a Fourier Quasicrystal on
The proof of this will occupy the remainder of this subsection. Define Then is a rational function holomorphic in A computation gives
therefore
where is the constant term in the Laurent expansion of the restriction
Express (149) in terms of the components of where for The chain rule gives the trigonometric series
when or equivalently is the constant term in this series.
Lemma 17
There exist and such that and are tempered Radon measures and
Proof. Since is a linear combination of and the coefficient of is the coefficient of in the trigonometric series expansion of and The frequencies appearing in this series are in the subgroup generated by This subgroup equals (149) implies
Since the coefficient of each exponential in this expansion is bounded by The set of frequencies of is finite and contained in the interior of a closed pointed cone (contains no line) defined by the vertex of the Newton polytope Therefore the set of frequencies of the expansion satisfies
A combinatorial argument shows that there exists such that
Multiplication by the trigonometric polynomial preserves the asymptotic polynomial bound on the distribution of frequencies. This concludes the proof.
Lemma 18
There exist and such that and are tempered Radon measures and
Proof. Follows from preceding results by defining
where are the combinatorial coefficients.
Lemma 19
There exists a discrete subset and a sequence such that and are tempered Radon measures and
Proof. As and the corresponding coefficients converge. Therefore is a Cauchy sequence in the Hausdorff metric for every so converges pointwise to a discrete set and the coefficients converge to so that defined above satisfy the asserted properties.
Corollary 5
5.4 Proof of Theorem 3
Assume that is real-rooted and uniformly generic. If is unfolded then Proposition 10 implies that is a Fourier quasicrystal on If is folded then for every and choose a random unit vector and then for every define be the trigonometric polynomial whose frequency replaced by frequency and whose amplitude at equals the amplitude of at Define Since foldedness is defined by equalities, with probability one the trigonometric map is unfolded. Therefore there exist unit vectors such that is unfolded for every . Then Proposition 10 implies is a Fourier quasicrystal on Since and are uniformly bounded for sufficiently small the following analogue of Proposition 8 holds.
converges to uniformly on bounded subsets of so the same argument used to prove Proposition 10 implies that is a Fourier quasicrystal.
6 Conclusions and Research Questions
Section 2 provided a background in harmonic and functional analysis relevant to almost periodic measures and Fourier quasicrystals. Proposition 2 implies that every Fourier quasicrystal on is a Bohr almost periodic measure. Example 2 constructs a set whose associated measure is a Bohr almost periodic Poisson measure with a dense spectrum. It raises the following question which was asked by Favorov in a personal communication:
Question 1
Do there exist multisets in whose associated measures are Bohr almost periodic but not Poisson?
If is a lattice subgroup and is Bohr almost periodic, then Bochner’s result (21) implies that the multiset parameterized by the function defined by is Bohr almost periodic, and hence its associated measure is Bohr almost periodic. Favorov ([22], Theorem 1) proved that all Bohr almost periodic multisubsets in arise in this way. The Fourier series of the continuous
does not converge absolutely. Choose and define by
Then is Bohr almost periodic and its spectrum is a dense subset of Moreover, its Fourier-Bohr transform is not even absolutely convergent when restricted to any nonempty open subset of This pathological behavior suggests:
Conjecture 1
The Bohr almost periodic measure associated to the multiset parameterized by the function is not a Poisson measure.
is of toral type with compactification hence its Fourier-Bohr transform has a expression derived in [50]. Therefore the
criteria for absolute convergence described by Zygmund [90] may be useful to answer this conjecture.
The Fourier quasicrystals constructed in Example 3 and in Section 5 have positive imteger weights and are of toral type.
This raises the following:
Question 2
Are all Fourier quasicrystals with positive imteger weights of toral type?
The result of Olevsly and Ulanovsky [72] implies that every Fourier quasicrystal on with positive imteger weights is of toral type. Moreover, we have been unable to construct Fourier quasicrystals on with positive imteger weights that are not of toral type.
Example 3 constructed Fourier quasicrystals using algebraic varieties that are generally not complete intersections, but we used them in Section 5
because Proposition
7 required a result in [10] about complete intersections. This raises:
Question 3
Can the sufficient conditions assumed in Section 5 be relaxed. Specifically, can the complete intersection assumption be removed?
Question 3 was answered affirmatively by Alon, Kummer, Kurasov and Vinzant [2]. They constructed Delone Fourier quasicrystals:
where is a strict Lee-Yang variety
([2], Definition 2.2) and
where
is a matrix all of whose minors are positive. If
with is a
Laurent map whose zero set equals then is -stable
by Definition 18 since the condition on defines an
open set of matrices. Therefore Lemma 13 implies that
the construction in [2] includes our constructions in Section 5,
but it is more general because it allows to be a non complete intersection.
Example 3 and [2] constructed irreducible Delone
sets whose associated measures are Fourier quasicrystals.
Question 4
Is there a method to construct irreducible non Delone Fourier quascrystals?
A proof of the following conjecture may be useful to construct non Delone Fourier quasicrystals of toral type.
Conjecture 2
If is an irreducible non Delone Bohr almost periodic multiset in of toral type that is associated with a compactification of then is the multipicity variety defined by a continous map
that is transverse to the foliation on and not injective. The transversality condition on implies that it induces an injection of fundamental groups
The method used in Section 5 in [50] to describe the Delone case must be refined to prove Conjecture 2.
Acknowledgments The authors thank Lior Alon, Sergey Favorov, Yves Meyer, Alexander Olevskii, and Peter Sarnak for sharing their knowledge about Fourier quasicrystals and Lang’s conjecture.
We thank the reviewer for suggesting revisions that improved the papers accuracy and clarity.